• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 08:00:08 on February 13, 2007

    Entry number 192865

    keyword=2/13 Owl Shift Summary

    00:00 Change target to C for production data on kinematics Y. Request 1uA but readjust current down to 0.5uA because of high rates. Need to set high pre-scale factors (PS3) to keep DAQ rate at about 1kHz.

    00:18 First production run on carbon for this kinematics.

    00:35 Lost communications with left arm. Reboot IOC. Inadvertantly over-rebooted.

    00:43 Restarted running on C target. The IOC & settings came back OK. Did NOT reset the p0 setpoint lest the magnets try to set themselves.

    00:59 Requested change to B4C target.

    01:05 Begin running on B4C target--Note Very High prescales, Alexander requested beam currents not drop below 0.5uA.

    01:39 Request target change to Li.

    01:45 Requested beam back from MCC.

    02:21 Request target change to Ta & set HRS-L angle to 14.0 deg

    02:26 Request 0.5uA from MCC, new kinematics X

    03:04 Request target change to C.

    03:09 Request 0.2uA from MCC because rates in T3 were 70kHz.

    03:15 Started run on carbon target with 0.2uA (tune beam!). I guess we'll see how this goes...

    03:17 MCC called to let us know that Hall C is moving the half wave plate from in to out.

    03:21 Presumably because the beam is in tune mode, we aren't getting beam position info. at these low currents.

    03:56 Request change to B4C target.

    04:00 Start running on B4C target. Large prescale (180) for T3, but the machine is going about as low as it will go!

    04:31 Request target change to Li.

    04:40 Resume production data taking on Li target.

    05:10 Request target change back to Ta & move spectrometer back to 16 deg.

    05:15 Back to kinematics Y, Ta target in, requested 0.5uA from MCC.

    05:09 Resume data taking on Ta in kinematics Y at 0.5uA

    5:50 Request target change to carbon target

    05:55 Resume data taking on C target.

    06:11 QE started (MCC didn't call!) No beam for a few minutes.

    06:22 Restarted running on carbon target

    06:37 Requested change of target to B4C.

    06:43 Restart data taking with B4C target

    07:13 Request change of target to Li. GUI gave its usual error.

    07:18 Request return of beam on Li target.

    07:21 Resume data taking on Li target.

    07:53 Request target change back to Ta. Set up for kinematics Z.

    07:59 Finished setting up for kinematics Z. Requested beam back.

    End of Shift!